OpenStudy (anonymous):

rewrite with positive exponents. assume that even roots are nonnegative quantities and that all are denominators are zero (4xy^-2)^-2

OpenStudy (anonymous):

im confused is the answer (4xy^-2)^-2 = 4xy^4 OR y^4/16*x^2

OpenStudy (lgbasallote):

\[\large \begin{array} )(4xy^{-2})^{-2} \implies \frac{1}{(4xy^{-2})^2}\implies \frac{1}{16x^2y^{-4}} \implies \frac{y^4}{16x^2}\end{array}\] does that help?

OpenStudy (anonymous):

yes it does thank you!

OpenStudy (lgbasallote):

welcome ^_^